OpenStudy (anonymous):

How do I find the general solution of dy/dx = 2/3 * (130 - y)?

OpenStudy (gorv):

bring terms having y on one side and x terms on other side

OpenStudy (gorv):

\[\frac{ dy }{ 130-y }=\frac{ 2 }{ 3 } *dx\]

OpenStudy (gorv):

now integrate it

OpenStudy (anonymous):

thanks!

OpenStudy (gorv):

\[\int\limits_{}^{} \frac{ dy }{ 130-y }=\frac{ 2 }{ 3 }* \int\limits_{}^{} dx\]